OpenStudy (anonymous):

can someone help me with a k12 math question on slopes please, I will give a medal Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3 A. y-7=2/3(x-1) B.y+7=2/3(x+1) C.y-7=-2/3(x+1) D.y+7=-2/3(x-1)

OpenStudy (anonymous):

I will fan and medal for help, I dont care about exact answers, I would just like to know how to do this

OpenStudy (anonymous):

is it b.? @confxsedaf I think...

OpenStudy (anonymous):

Yes it is b

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

will you explain why please, you dont have to though...

OpenStudy (anonymous):

No I am not I am only a senior???

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I said no I do not understand it is just that I am in 12th grade

OpenStudy (anonymous):

im in 8th XD

OpenStudy (anonymous):

I think its b...

OpenStudy (anonymous):

YEES it is b!

OpenStudy (anonymous):

ok...I will tell you if it was in a sec.

OpenStudy (confxsedaf):

no its d

OpenStudy (confxsedaf):

i think sorry for not reply soon

OpenStudy (anonymous):

it was d...DANG IT! but anyway heres a medal for helping XD

OpenStudy (confxsedaf):

ah sorry! i was held up cuz i kept getting -19/3 which was the points hope you know how to do it now

OpenStudy (confxsedaf):

here is a website that teaches you how to do it i used it when i was in 8th grade http://www.webmath.com/equline2.html

OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2? a. y+3=2(x+5) b. y-3=2(x-5) --> c. y-5=2(x-3) d. y-5=-2(x-3) C. is the answer

OpenStudy (confxsedaf):

yes that one is C'

OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.55 2.25 2.95 3.65 Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.014x + 49.15 B. y = 0.014x + 1.55 -->C. y = 0.014x + 0.85 D. y = 58.8x + 0.85 the answer is c.

OpenStudy (anonymous):

actually I was putting what i got, I got an 80% and these were the correct answers

OpenStudy (confxsedaf):

yep 1/3 is the slope for that graph ones

OpenStudy (confxsedaf):

what branch in k12 do u go to? i'm in keystone

OpenStudy (anonymous):

cava

OpenStudy (anonymous):

im also in 8th grade

OpenStudy (confxsedaf):

im in 10 did you understand how to get slopes now

OpenStudy (anonymous):

yep, im gonna close this question now XD

OpenStudy (confxsedaf):

kk

OpenStudy (anonymous):

other people that have the same questions as these, this will get them a 100% XD